your job requires you to cut 4 sides of an 86 x 28 rectangle how many total feet will you cut?
Answer by stanbon(75887)

your job requires you to cut 4 sides of an 86 x 28 rectangle how many total feet will you cut?
-----------
perimeter = 2(86+28) = 2(114) = 228 ft.
==========================================
Cheers,
Stan H.